OpenStudy (anonymous):

The length of a rectangular garden is 6m greater than the width. The area is 72m^2. Find the dimensions of the garden. The width of the garden is The length of the garden is

OpenStudy (saifoo.khan):

Area = l*b 72 = (x+6)(x) Solving so x, We will get, 6 and -12.

OpenStudy (saifoo.khan):

So, Length = 12 width = 6

OpenStudy (anonymous):

Thanks

OpenStudy (saifoo.khan):

Welcome.